40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 22 23:53:16 2019, gautam, Update, ALS, IR ALS locking attempt ALS_X_outOfLoopnoise.pdfALS_X_inLoopnoise.pdfCARM_loopShape.pdf
    Reply  Wed Jul 24 00:22:24 2019, gautam, Update, ALS, PSL frequency locked to XARM length using ALS ALS_X_noise_POX.pdfBR_comparison.pdfALS_CARM_OLG.pdf
       Reply  Thu Jul 25 12:25:56 2019, gautam, Update, ALS, IR ALSX noise ALS_broadPeaks.pdf
Message ID: 14811     Entry time: Thu Jul 25 12:25:56 2019     In reply to: 14802
Author: gautam 
Type: Update 
Category: ALS 
Subject: IR ALSX noise 

Summary:

  1. There are some broad peaks in the ALS out-of-loop noise, centered at ~145 Hz, ~245 Hz and ~570 Hz which are absent in both the POX in-lock error signal and in the green PDH error signals (see Attachment #1). So I conclude they originate in the IR ALS beat chain somewhere. Needs more investigation, in the general quest to improve the ALS noise.
  2. This measurement also shows that the ALS noise is limited by unsuppressed EX green PDH frequency noise above ~400 Hz (100 Hz if you ignore the unexplained broad humps).

These spectra were taken with the arm cavity length locked to the PSL frequency using POX as an error signal, and the EX laser frequency locked to the XARM cavity length by the analog PDH servo at EX, so there is no feedback control with the ALS beat signal as an error signal.

Other details:

  • The transition of arm resonance control from POX to ALS error signal is more robust now - I am able to do this during daytime, and also maintain the lock for >20 minutes at a time.
  • Rana encouraged me not to spend too much time on this - so my next goal here will be to get the Y arm IR ALS working, and then we can control the two arms using ALS error signals in the CARM/DARM basis instead of the X/Y basis.
  • I still think it's worth getting the ALS good enough that the locking becomes repeatable and reliable.
    • The main task here is going to be re-doing the EY green layout to match the EX layout, get good MM into the cavity etc.
    • The IR light also has to be coupled into the fiber at EY.
Attachment 1: ALS_broadPeaks.pdf  581 kB  Uploaded Thu Jul 25 13:47:00 2019  | Hide | Hide all
ALS_broadPeaks.pdf
ELOG V3.1.3-